OpenStudy (anonymous):

Which is the solution of the system? A. (-4,2) B. (-3,1) C. (-3,-5) D. (4,-6)

OpenStudy (anonymous):



OpenStudy (anonymous):

do i use distributive property???

OpenStudy (astrophysics):

I don't really know what your question is asking

OpenStudy (anonymous):

yeah nvm i understand it now that i really look at it..its all good

OpenStudy (astrophysics):

\[\frac{ (x+y) }{ (2x-3y) } = -\frac{ 2 }{ 9 }\] is the problem this?

OpenStudy (astrophysics):

Can you just take a picture of it and upload it

OpenStudy (astrophysics):

Oh, that makes sense, they are two different equations right?

OpenStudy (astrophysics):

Yeah!

OpenStudy (anonymous):

so basically i just use distributive property?

OpenStudy (astrophysics):

Ok so you have two equations, \[x+y=-2\] \[2x-3y=-9\] No not quite, so does it ask for any methods, like elimination, substitution, matrices?

OpenStudy (anonymous):

it asks what the solution is

OpenStudy (astrophysics):

Ok, we can use any method then

OpenStudy (anonymous):

im gonna give you a medal just cuz youve been the most help ive had since ive gotten on here

OpenStudy (astrophysics):

Lets use substitution, so to do this problem, we want to isolate one of the variables in one of the equations and plug it into the second equation, that will help us find the x value and then we can go back and find the y value. We could also take your options in to our advantage and do it the lazy way but I don't think you will learn that way.

OpenStudy (astrophysics):

Ok I think drawing it out will make more sense, first you want to label the equations

OpenStudy (anonymous):

Did you read what i said?

OpenStudy (astrophysics):

 Hey, yes I did thanks :) So lets do this problem now, pick one of the equations and one of the variables to isolate in that equation.